# Situationally-Aware Dynamics Learning

### v2

@@ Abstract
Autonomous robots operating in complex, unstructured environments face significant challenges due to latent, unobserved factors that obscure their understanding of both their internal state and the external world. Addressing this challenge would enable robots to develop a more profound grasp of their operational context. To tackle this, we propose a novel framework for online learning of hidden state representations, with which the robots can adapt in real-time to uncertain and dynamic conditions that would otherwise be ambiguous and result in suboptimal or erroneous behaviors. Our approach is formalized as a Generalized Hidden Parameter Markov Decision Process, which explicitly models the influence of unobserved parameters on both transition dynamics and reward structures. Our core innovation lies in learning online the joint distribution of state transitions, which serves as an expressive representation of latent ego- and environmental-factors. This probabilistic approach supports the identification and adaptation to different operational situations, improving robustness and safety. Through a multivariate extension of Bayesian Online Changepoint Detection, our method segments changes in the underlying data generating process governing the robot’s dynamics. The robot’s transition model is then informed with a symbolic representation of the current situation derived from the joint distribution of latest state transitions, enabling adaptive and context-aware decision-making. To showcase the real-world effectiveness, we validate our approach in the challenging task of unstructured terrain navigation, where unmodeled and unmeasured terrain characteristics can significantly impact the robot’s motion. Extensive experiments in both simulation and real world reveal significant improvements in data efficiency, policy performance, and the emergence of safer, adaptive navigation strategies. Accompanying video: https://youtu.be/VKR18WaSCAk.
I Introduction
@@
Figure 1: Motivating Example. When the robot has access only to incomplete state information, the same action, $\mathbf{a}$, can yield significantly different outcomes. In this example, unmodeled terrain types—factors absent from the state and transition model—affect the robot’s response, resulting in distinct changes in position and heading ($\Delta x$, $\Delta y$, and $\Delta\psi$, respectively) for the same control action (commanded linear and angular velocity). Generalized Hidden Parameter Markov Decision Processes (GHP-MDPs) address this challenge by modeling the unknown latent factors responsible for these discrepancies and incorporating them into the transition model, thereby enhancing predictive accuracy. (Illustrative environment images generated with Gemini AI.)
@@
For reliable robot deployment in unstructured and dynamic environments, it is crucial for robotic systems to adapt to unforeseen challenges and operate effectively under uncertainty. Consider, for example, an Unmanned Ground Vehicle (UGV) navigating steep and rugged terrain. Such a robot may become immobilized due to unexpected factors, ranging from unmodeled terrain properties to variations in its own internal dynamics. These factors are inherently difficult, if not impossible, to anticipate exhaustively [du2011robot, kurniawati2012global, green2020planning, siva2021enhancing, yin2023reliable, xu2024boundary]. This challenge is further exacerbated by the impracticality of equipping robots with every conceivable sensor and analysis algorithm needed to fully capture the world’s and robot’s full state in all scenarios. Consequently, robots must learn to operate effectively using only partial information derived from their observable state—the limited set of measurements and estimates available in real time. This necessitates the development of adaptive mechanisms capable of inferring and representing the hidden aspects of the robot and its environment to inform decision-making and enhance resilience in uncertain conditions [mccallum1996hidden, ferguson2004pao, kroemer2014learning, kobayashi2022latent].
@@
Furthermore, when a robot relies on partial information to operate on a complex environment, the risk of ambiguous scenario representation becomes significant [meyer2011using, pramanick2022talk, celemin2023knowledge]. Such ambiguity is undesirable, as it increases uncertainty during decision-making, often leading to suboptimal plans and, in some cases, unsafe or hazardous behaviors. This challenge is exemplified in Figure 1: a UGV with incomplete state information—lacking a (robust) scene interpretation module to detect the terrain type it is traversing—executes an action from what appears to be the same state. However, the outcomes of these actions differ drastically, revealing a discrepancy that cannot be attributed solely to aleatoric errors (those arising from process or observation noise). This highlights the critical need for robust methods to infer and resolve hidden state ambiguities to ensure reliable robot operation.
@@
To address this, we propose a framework for learning online a representation of the unmodeled and unobserved latent factors that cause such ambiguities. This task is challenging because it requires balancing representation capacity, data- and learning-efficiency, and minimizing inductive biases that might inadvertently exclude critical latent factors [langdon2019uncovering, hundt2006representing]. Motivated by these constraints, we introduce an online hidden-state representation learning method that enables the robot to recover and leverage compact representations of latent world and robot factors in real time. By exploiting this learned representation the robot can dynamically adapt its behavior and reliably accomplish tasks even in uncertain and unstructured environments. Prior work has attempted to address this by learning distributions over causal structures—an approach explicitly intended to account for uncertainty in the dynamics model caused in part by unobserved latent factors [murillogonzalez2025causaldistributions]; our method instead targets compact, real-time latent state representations for online adaptation.
@@
Recent studies have introduced various methods that leverage learned auxiliary representations during dynamics or policy learning to enable robot adaptation [lee2020learning, kumar2021rma, kumar2022adapting, qi2023hand, liang2024rapid]. However, these methods are data-intensive and often rely on multi-stage procedures that require privileged information, such as ground-truth data that is difficult to obtain or is only available in simulation. To overcome these limitations, our proposed method learns such representations in an online and unsupervised manner, thus eliminating the need for multi-stage approaches. We achieved this by formulating the problem within the framework of Generalized Hidden Parameter Markov Decision Processes (GHP-MDPs) [perez2020generalized, doshi2016hidden].
@@
An important component of our framework is to represent the hidden state by learning online the joint distribution of the state transitions. We refer to this state transition distribution at a given moment as the situation of the robot. This approach enables the robot to capture an expressive probabilistic representation of its current context, encompassing both its state and the underlying environmental conditions. By leveraging proprioceptive sensing, such as inertial measurements, to assess the robot’s state, this representation becomes sufficiently general to effectively disambiguate the robot’s state across various scenarios. Additionally, as illustrated in Figure 1, the state evolution exhibits diverse probability density profiles under different conditions. This diversity facilitates the identification of latent factors that govern distinct scenarios. In this way, the situation can be understood as the Underlying Data Generating Process (UDGP)—a conceptual model that characterizes the true, often hidden, factors governing the system’s dynamics at any given moment [Goodfellow-et-al-2016, p. 109]. The UDGP encapsulates the combined influence of both robot-specific and environmental variables that affect the robot’s behavior, enabling a more accurate understanding of its state. To obtain this situation distribution in real-time as the robot operates, we introduce a multivariate extension of Bayesian Online Changepoint Detection (BOCD) [adams2007bayesian]. This extension allows us to model the situation the robot is currently operating under and detect when a situation change occurs.
@@
  • To validate our framework, we applied our method to control a UGV navigating unstructured terrain, where real-time adaptation and online motion planning are essential for overcoming unexpected challenges and increasing the likelihood of successful traversal. We demonstrate that learning our situationally-aware dynamics model using Model-Based Reinforcement Learning (MBRL) is significantly more data-efficient and results in superior policies compared to competitive reinforcement learning (RL) and physics-based approaches. Additionally, we observe the emergence of adaptive navigation behaviors that enhance the safety and effectiveness of the UGV’s traversal through challenging environments, facilitated by the augmented situational awareness provided by our hidden state representation.
@@
The remainder of this work is structured as follows. Section II reviews the relevant literature, focusing on policy and dynamics learning methods that incorporate learned auxiliary representations. Section III introduces GHP-MDPs and compares them to related MDP-derived families. We also present the problem of changepoint detection. In Section IV, we detail our proposed approach, describing how to model transition dynamics online, represent latent factors, and integrate them into the transition model in a data-efficient way. In Section V, we demonstrate that our method yields improved and safer adaptation in both simulation and real-world experiments. Finally, Sections VI and VII discuss our findings and outline the conclusions drawn from our work, respectively.
II Related Work
@@
Our work focuses on efficiently learning a robot’s dynamics model in conjunction with a hidden state representation to enable safe and effective adaptation in ambiguous scenarios. We review relevant literature on planning and learning with auxiliary representations—both latent and symbolic—along with research that specifically addresses the problem of unstructured terrain navigation.
II-A Neural-Symbolic Learning in Robotics
@@
The field of neural-symbolic learning consists of two primary branches [wang2024imperative]. We will first focus on methods rooted in discrete logical reasoning and guiding abstractions. The second branch involves Physics-Infused Neural Networks (PINNs), which handle continuous abstractions to represent physical characteristics, which we cover later on.
@@
Analogous to our work, gopalan2020simultaneously used changepoint detection for robotic applications. They use it to segment demonstration trajectories from which the robot learns skills. Then they learn symbols by identifying the termination conditions of the skills via clustering. Their objective was to map behaviors to natural language commands. However, for training they rely on demonstration trajectories that have been paired to natural language instructions, and find the symbols from the provided dataset. Conversely, in our case the agent discovers the symbols online and unsupervisedly, from the data generated by its proprioception, while it is exploring the environment.
@@
[konidaris2012robot] construct skill trees from demonstration trajectories. Each skill has a goal and an abstraction and they can later be composed. Similarly, they use a changepoint detection algorithm in charge of determining when the most relevant abstraction changes. However, they cannot recognize repeated skills and require the availability of an abstraction library while we directly model the robot’s dynamics which can then be used with sampling-based planners. Other authors have also worked along this line [kulic2009online, chiappa2010movement, niekum2012learning]. Moreover, learned symbolic abstractions for planning, skill learning and control have been explored in [silver2022learning, achterhold2023learning, belta2007symbolic, kumar2023learning, rosen2023synthesizing].
@@
Research that is in proximity also includes ly2012learning which propose a method for learning a system’s hybrid dynamics directly from time-series data through multi-modal symbolic regression. Similarly, [kim2022physics] focus on estimating key physical parameters, such as tire slippage and longitudinal force, in the context of an autonomous vehicle model, offering critical insights for improving vehicle performance and control. Closer to our work, [zhao2024physord] employ physics-informed neural networks (PINNs) to predict motion in off-road driving scenarios, addressing the challenges posed by rough and unpredictable terrain while ensuring more accurate trajectory forecasting.
II-B Model-Based Reinforcement Learning
@@
System Identification (SI) aims to discover and model the relationship between the inputs and outputs of a system [aastrom1971system, ljung2010perspectives]. MBRL is a SI method based on balancing exploration and exploitation. It iteratively collects more informative data for the task at hand, thanks to planning with an improved model learned with the data collected in previous iterations. Thus, the model evolves, leading to better planning performance which results in exploring more informative sections of the state-action space [sutton2018reinforcement].
[abbeel2006using] explore how to use an approximate model with few real-life trials. They evaluate a policy using a sample trial from the real world, and leverage the approximate model to estimate the derivative of the evaluation with respect to the policy parameters. [deisenroth2011pilco] introduce a model-based policy search strategy for high-dimensional control tasks. They employ Gaussian Processes to learn the dynamics and incorporate the model’s uncertainty during planning and policy evaluation. They highlight the data efficiency of this approach. chua2018deep propose PETS, which uses ensembles of uncertainty-aware neural-networks together with sampling-based optimizers for planning.
@@
hansen2023td presented TD-MPC2, which learns five components consisting of an (i) observation encoder, (ii) latent-space dynamics model, (iii) transition reward predictor, (iv) return or terminal value predictor, and (v) a policy predicting return-maximizing actions; to predict returns from a sequence of actions, usually from high-dimensional observations. Other works have also learned latent dynamics models [ha2018world, hafner2019dream]. MoDem-V2 [lancaster2023modem] builds on top of TD-MPC to learn vision-based models in the real-world for robot manipulation tasks. They identify unsafe exploration and over-optimism as a significant failure point for applying MBRL in the real-world and develop three techniques for safer real-world learning: policy centering, agency transfer and actor-critic ensembles.
II-C Policy and Dynamics Learning with Learned Auxiliary Representations
@@
II-C1 Model-Free Reinforcement Learning:
[lee2020learning] use a temporal convolutional network that takes a window of past proprioceptive states to generate control commands for quadruped locomotion over challenging terrains. This model is trained to “implicitly reason about contact and slippage events” from the proprioception. However, learning the previous model requires a teacher policy that has been previously obtained using privileged information of the terrain and the robot’s contact with it. Furthermore, they also make use of an adaptive curriculum in charge of modulating the terrain’s difficulty according the current performance of the policy. [kumar2021rma] proposed Rapid Motor Adaptation (RMA), which is a two-stage process to train adaptive policies via Model-Free RL (MFRL). First, an “environmental factors encoder” model is trained to map “privileged information” only available in simulation to a latent vector. This latent vector is used as input together with the current state and action to a “base-policy” trained via Proximal Policy Optimization [schulman2017proximal, PPO]. Then, during the second phase, an “adaptation module” is trained to map a fixed-size window of past state-action pairs into a latent vector optimized to match the latent representations produced by the “environmental factors encoder”. This way, during deployment, the latent representation can be obtained without the need for privileged information. [kumar2022adapting] add a third phase to the RMA framework to enable adaptive bipedal locomotion. The new stage involves using PPO to fine-tune the base policy but maintaining the adaptation module weights frozen, to ensure that the policy supports imperfect estimation of the latent representation. Follow up works to RMA include [qi2023hand] which also uses proprioception history and privileged information to train a model for in-hand object rotation. Also, [liang2024rapid] learn the latent representation from depth perception for manipulation tasks.

### v3

# Lantao Liu

@@ Abstract
Autonomous robots operating in complex, unstructured environments face significant challenges due to latent, unobserved factors that obscure their understanding of both their internal state and the external world. Addressing this challenge would enable robots to develop a more profound grasp of their operational context. To tackle this, we propose a novel framework for online learning of hidden state representations, with which the robots can adapt in real-time to uncertain and dynamic conditions that would otherwise be ambiguous and result in suboptimal or erroneous behaviors. Our approach is formalized as a Generalized Hidden Parameter Markov Decision Process, which explicitly models the influence of unobserved parameters on both transition dynamics and reward structures. Our core innovation lies in learning online the joint distribution of state transitions, which serves as an expressive representation of latent ego- and environmental-factors. This probabilistic approach supports the identification and adaptation to different operational situations, improving robustness and safety. Through a multivariate extension of Bayesian Online Changepoint Detection, our method segments changes in the underlying data generating process governing the robot’s dynamics. The robot’s transition model is then informed with a symbolic representation of the current situation derived from the joint distribution of latest state transitions, enabling adaptive and context-aware decision-making. To demonstrate effectiveness, we validate our approach on an unmanned ground vehicle operating in diverse unstructured terrains, both in simulation and in real-world experiments. We also evaluate a quadrotor in simulation under randomly changing wind conditions. Both setups introduce unmodeled and unmeasured environmental factors that substantially affect robot motion. Extensive experiments in both simulation and real world reveal significant improvements in data efficiency, policy performance, and the emergence of safer, adaptive navigation strategies. Website: https://alejandromllo.github.io/research/situational-awareness/.
keywords:
Hidden State Representation, Symbolic Reasoning, Representation Learning, Online Learning, MBRL
1 Introduction
@@
Figure 1: Motivating Example. When the robot has access only to incomplete state information, the same action, $\mathbf{a}$, can yield significantly different outcomes. In this example, unmodeled terrain types—factors absent from the state and transition model—have an effect on the robot’s dynamics, resulting in distinct changes in position and heading ($\Delta x$, $\Delta y$, and $\Delta\psi$, respectively) for the same control action (commanded linear and angular velocity). Generalized Hidden Parameter Markov Decision Processes (GHP-MDPs) address this challenge by modeling the unknown latent factors responsible for these discrepancies and incorporating them into the transition model, thereby enhancing predictive accuracy. (Illustrative environment images generated with Gemini AI.)
@@
For reliable robot deployment in unstructured and dynamic environments, it is crucial for robotic systems to adapt to unforeseen challenges and operate effectively under uncertainty. Consider, for example, an Unmanned Ground Vehicle (UGV) navigating steep and rugged terrain. Such a robot may become immobilized due to unexpected factors, ranging from unmodeled terrain properties to variations in its own internal dynamics. These factors are inherently difficult, if not impossible, to anticipate exhaustively Du Toit and Burdick (2011); Kurniawati et al. (2012); Green et al. (2020); Siva et al. (2021); Yin et al. (2023); Xu et al. (2024b). This challenge is further exacerbated by the impracticality of equipping robots with every conceivable sensor and analysis algorithm needed to fully capture the world’s and robot’s full state in all scenarios. Consequently, robots should learn to operate effectively using only partial information derived from their observable state—the limited set of measurements and estimates available in real time. This necessitates the development of adaptive mechanisms capable of inferring and representing the hidden aspects of the robot and its environment to inform decision-making and enhance resilience in uncertain conditions McCallum (1996); Ferguson et al. (2004); Kroemer et al. (2014); Kobayashi et al. (2022).
@@
Furthermore, when a robot relies on partial information to operate in a complex environment, the risk of ambiguous scenario representation becomes significant Meyer-Delius et al. (2011); Pramanick et al. (2022); Celemin and Kober (2023). Such ambiguity is undesirable because it increases uncertainty during decision-making, often producing suboptimal plans or, in extreme cases, unsafe behavior. This challenge is illustrated in Figure 1: a UGV with incomplete state information—lacking a (robust) scene interpretation module to detect the terrain it is traversing—executes the same action from what appears to be an identical initial state, yet the observed outcomes differ dramatically; these discrepancies cannot be explained solely by aleatoric noise (process or observation noise). While we use terrain as an easy-to-visualize example, many other unobserved factors can similarly perturb dynamics (e.g., payload shifts, wheel or tire wear, actuator degradation, sensor bias, unseen vegetation or obstacles, local slope/contact conditions, etc.). These sources often introduce systematic, context-dependent shifts that complicate learning and control, underscoring the need for methods that can infer and resolve hidden-state ambiguities to ensure reliable robot operation.
@@
To address this, we propose a framework for learning online a representation of the unmodeled and unobserved latent factors that cause such ambiguities. This task is challenging because it requires balancing representation capacity, data- and learning-efficiency, and minimizing inductive biases that might inadvertently exclude critical latent factors Langdon et al. (2019); Hundt et al. (2006). Motivated by these constraints, we introduce an online hidden-state representation learning method that enables the robot to recover and leverage compact representations of latent world and robot factors in real time. By exploiting this learned representation the robot can dynamically adapt its behavior and reliably accomplish tasks even in uncertain and unstructured environments. Prior work has attempted to address this by learning distributions over causal structures—an approach explicitly intended to account for uncertainty in the dynamics model caused in part by unobserved latent factors Murillo-González et al. (2025); our method instead targets compact, real-time latent state representations for online adaptation.
@@
Recent studies have introduced various methods that leverage learned auxiliary representations during dynamics or policy learning to enable robot adaptation Lee et al. (2020a); Kumar et al. (2021, 2022); Qi et al. (2023); Liang et al. (2024). However, these methods are data-intensive and often rely on multi-stage procedures that require privileged information, such as ground-truth data that is difficult to obtain or is only available in simulation. To overcome these limitations, our proposed method learns such representations in an online and unsupervised manner, thus eliminating the need for multi-stage approaches. We achieved this by formulating the problem within the framework of Generalized Hidden Parameter Markov Decision Processes (GHP-MDPs) Perez et al. (2020); Doshi-Velez and Konidaris (2016).
@@
An important component of our framework is to represent the hidden state by learning online the joint distribution of the state transitions. We refer to this state transition distribution at a given moment as the situation of the robot. This approach enables the robot to capture an expressive probabilistic representation of its current context, encompassing both its state and the underlying environmental conditions. By leveraging proprioceptive sensing, such as inertial measurements, to assess the robot’s state, this representation becomes sufficiently general to effectively disambiguate the robot’s state across various scenarios. Additionally, as illustrated in Figure 1, the state evolution exhibits diverse probability density profiles under different conditions. This diversity facilitates the identification of latent factors that govern distinct scenarios. In this way, the situation can be understood as the Underlying Data Generating Process (UDGP)—a conceptual model that characterizes the true, often hidden, factors governing the system’s dynamics at any given moment (Goodfellow et al., 2016, p. 109). The UDGP encapsulates the combined influence of both robot-specific and environmental variables that affect the robot’s behavior, enabling a more accurate understanding of its state. To obtain this situation distribution in real-time as the robot operates, we introduce a multivariate extension of Bayesian Online Changepoint Detection (BOCD) Adams and MacKay (2007). This extension allows us to model the situation the robot is currently operating under and detect when a situation change occurs.
@@
  • We validate the framework on a UGV navigating unstructured terrain (simulation and real world), where the dynamics model is learned via active data collection within a Model-Based Reinforcement Learning (MBRL) loop, and on a simulated quadrotor under changing winds, whose model is trained using a dataset of state transitions. This allows us to showcase the suitability of our method for both active and passive learning settings, as well as its applicability across different platform types. We observe that the situationally-aware models show substantially improved data efficiency and outperform RL baselines and physics-based controllers. Furthermore, the learned hidden-state representation facilitates the emergence of adaptive behaviors that increase the safety and effectiveness of the downstream planners.
@@
The remainder of this work is structured as follows. Section 2 reviews the relevant literature, focusing on policy and dynamics learning methods that incorporate learned auxiliary representations. Section 3 introduces GHP-MDPs and compares them to related MDP families. We also present the problem of changepoint detection. In Section 4, we detail our proposed approach, describing how to model transition dynamics online, represent latent factors, and integrate them into the transition model in a data-efficient way. In Section 5 we demonstrate that our approach yields safer and more data-efficient adaptation for a UGV, validated in both simulation and real-world trials. Section 6 presents complementary results for a quadrotor evaluated in simulation under randomly varying wind fields. Together, these evaluations show that our situationally-aware dynamics representation enables adaptive and more robust control across both ground and aerial platforms. Finally, Sections 7 and 8 discuss our findings and outline the conclusions drawn from our work, respectively.
2 Related Work
@@
Our work focuses on efficiently learning a robot’s dynamics model in conjunction with a hidden state representation to enable safe and effective adaptation in ambiguous scenarios. We review relevant literature on planning and learning with auxiliary representations—both latent and symbolic—along with research that specifically addresses the problem of robot motion control in unstructured environments.
2.1 Neural-Symbolic Learning in Robotics
@@
The field of neural-symbolic learning consists of two primary branches Wang et al. (2024). We will first focus on methods rooted in discrete logical reasoning and guiding abstractions. The second branch involves Physics-Infused Neural Networks (PINNs), which handle continuous abstractions to represent physical characteristics, which we cover later on.
@@
Analogous to our work, Gopalan et al. (2020) used changepoint detection for robotic applications. They use it to segment demonstration trajectories from which the robot learns skills. Then they learn symbols by identifying the termination conditions of the skills via clustering. Their objective was to map behaviors to natural language commands. However, for training they rely on demonstration trajectories that have been paired to natural language instructions, and find the symbols from the provided dataset. Conversely, in our case the agent discovers the symbols online and unsupervised, from the data generated by its proprioception, while it is exploring the environment.
@@
Konidaris et al. (2012) construct skill trees from demonstration trajectories. Each skill has a goal and an abstraction and they can later be composed. Similarly, they use a changepoint detection algorithm in charge of determining when the most relevant abstraction changes. However, they cannot recognize repeated skills and require the availability of an abstraction library while we directly model the robot’s dynamics which can then be used with sampling-based planners. Other authors have also worked along this line Kulic et al. (2009); Chiappa and Peters (2010); Niekum et al. (2012). Moreover, learned symbolic abstractions for planning, skill learning and control have been explored in Silver et al. (2022); Achterhold et al. (2023); Belta et al. (2007); Kumar et al. (2023); Rosen et al. (2023).
@@
Research that is in proximity also includes Ly and Lipson (2012) which propose a method for learning a system’s hybrid dynamics directly from time-series data through multi-modal symbolic regression. Similarly, Kim et al. (2022) focus on estimating key physical parameters, such as tire slippage and longitudinal force, in the context of an autonomous vehicle model, offering critical insights for improving vehicle performance and control. Closer to our work, Zhao et al. (2024) employ physics-informed neural networks (PINNs) to predict motion in off-road driving scenarios, addressing the challenges posed by rough and unpredictable terrain while ensuring more accurate trajectory forecasting.
2.2 Model-Based Reinforcement Learning
@@
System Identification (SI) aims to discover and model the relationship between the inputs and outputs of a system Åström and Eykhoff (1971); Ljung (2010). MBRL is a SI method based on balancing exploration and exploitation. It iteratively collects more informative data for the task at hand, thanks to planning with an improved model learned with the data collected in previous iterations. Thus, the model evolves, leading to better planning performance which results in exploring more informative sections of the state-action space Sutton and Barto (2018).
@@
Abbeel et al. (2006) explore how to use an approximate model with few real-life trials. They evaluate a policy using a sample trial from the real world, and leverage the approximate model to estimate the derivative of the evaluation with respect to the policy parameters. Deisenroth and Rasmussen (2011) introduce a model-based policy search strategy for high-dimensional control tasks. They employ Gaussian Processes to learn the dynamics and incorporate the model’s uncertainty during planning and policy evaluation. They highlight the data efficiency of this approach. Chua et al. (2018) propose PETS, which uses ensembles of uncertainty-aware neural-networks together with sampling-based optimizers for planning. Hansen et al. (2024) presented TD-MPC2, which learns five components consisting of an (i) observation encoder, (ii) latent-space dynamics model, (iii) transition reward predictor, (iv) return or terminal value predictor, and (v) a policy predicting return-maximizing actions; to predict returns from a sequence of actions, usually from high-dimensional observations. Other works have also learned latent dynamics models Ha and Schmidhuber (2018); Hafner et al. (2019). MoDem-V2 Lancaster et al. (2023) builds on top of TD-MPC to learn vision-based models in the real-world for robot manipulation tasks. They identify unsafe exploration and over-optimism as a significant failure point for applying MBRL in the real-world and develop three techniques for safer real-world learning: policy centering, agency transfer and actor-critic ensembles.
2.3 Policy and Dynamics Learning with Learned Auxiliary Representations
@@
2.3.1 Model-Free Reinforcement Learning:
Lee et al. (2020a) use a temporal convolutional network that takes a window of past proprioceptive states to generate control commands for quadruped locomotion over challenging terrains. This model is trained to “implicitly reason about contact and slippage events” from the proprioception. However, learning the previous model requires a teacher policy that has been previously obtained using privileged information of the terrain and the robot’s contact with it. Furthermore, they also make use of an adaptive curriculum in charge of modulating the terrain’s difficulty according to the current performance of the policy. Kumar et al. (2021) proposed Rapid Motor Adaptation (RMA), which is a two-stage process to train adaptive policies via Model-Free RL (MFRL). First, an “environmental factors encoder” model is trained to map “privileged information” only available in simulation to a latent vector. This latent vector is used as input together with the current state and action to a “base-policy” trained via Proximal Policy Optimization (Schulman et al., 2017, PPO). Then, during the second phase, an “adaptation module” is trained to map a fixed-size window of past state-action pairs into a latent vector optimized to match the latent representations produced by the “environmental factors encoder”. This way, during deployment, the latent representation can be obtained without the need for privileged information. Kumar et al. (2022) add a third phase to the RMA framework to enable adaptive bipedal locomotion. The new stage involves using PPO to fine-tune the base policy but maintaining the adaptation module weights frozen, to ensure that the policy supports imperfect estimation of the latent representation. Follow up works to RMA include Qi et al. (2023) which also uses proprioception history and privileged information to train a model for in-hand object rotation. Also, Liang et al. (2024) learn the latent representation from depth perception for manipulation tasks.